# The externally definable Ramsey property and fixed points on type spaces

### v3

@@ 6.3. Free amalgamation classes via Nešetřil-Rödl
Let $n\geq 2$, and let =$\mathcal{L}=\{R_{i}:i\in I\}$ with each $R_{i}$ of arity $n$. Let $M$ be a Fraïssé $\mathcal{L}$-structure with age $\mathcal{K}$ such that each $R_{i}^{M}$ is symmetric and $\mathcal{K}$ has free amalgamation. Then $M$ has the externally definable Ramsey property.
@@
(Similar to [16, Example 5.6].) Let $\mathcal{K}_{\leq}$ be the order expansion of $\mathcal{K}$ with Fraïssé limit $M_{\leq}$ (with $M$ the $\mathcal{L}$-reduct of $M_{\leq}$). By Theorem 6.5), we see that $M_{\leq}$ has the Ramsey property. To show that $M$ has externally definable Ramsey, by Proposition 3.13 we need only check for $|\bar{a}|<n$ and atomic formulae. Let $i\in I$ and $\bar{a}\in\mathcal{K}$, $|\bar{a}|<n$. Let $\chi:\binom{M}{\bar{a}}\to 2$ be externally defined by $\bar{e}\in N\succeq M$ (where $|\bar{e}|+|\bar{a}|=n$) and $\phi(\bar{x},\bar{y})\equiv R_{i}(\bar{x},\bar{y})$. Let $B\in\mathcal{K}$.
@@ 6.4. The directed graphs $S(2)$ and $S(3)$
The dense local order $S(2)$ is defined as follows. Let =$\mathcal{L}=\{R\}$ with $R$ binary. Let $\mathbb{T}$ denote the unit circle in the complex plane, and let $S(2)$ be the $\mathcal{L}$-structure whose domain consists of the points of $\mathbb{T}$ with rational argument, and where $aRb$ iff the angle subtended at the origin by the directed arc from $a$ to $b$ is less than $\pi$. As the domain of $S(2)$ excludes antipodal points, $S(2)$ is a tournament. We now state a number of facts about $S(2)$.
We call a tournament $(A,R^{A})$ linear if $R^{A}$ is a linear order. Let $\overrightarrow{C}_{3}$ denote the directed $3$-cycle. It is easy to see that a tournament is linear iff it does not contain $\overrightarrow{C}_{3}$ as a subdigraph. Let $D^{+}_{4}$ denote the tournament on $4$ vertices given by a vertex with $3$ out-edges to $\overrightarrow{C}_{3}$, and let $D^{-}_{4}$ denote the tournament on $4$ vertices given by a vertex with $3$ in-edges from $\overrightarrow{C}_{3}$. We call a tournament $A$ a local order if, for each vertex $a\in A$, both the out-neighbourhood of $a$ and the in-neighbourhood of $a$ are linear tournaments. It is again straightforward to see that $A$ is a local order iff $A$ excludes $D^{+}_{4}$ and $D^{-}_{4}$ as subdigraphs.
We have that $S(2)$ is ultrahomogeneous, and its age is the class of all finite local orders. This is not entirely straightforward - see [18] for further details.
@@
Take $e\in S(2)$, and let $\phi(x,y)\equiv xRy$. Let $\chi$ be the vertex-colouring of $S(2)$ externally defined by $e\in S(2)$ and $\phi(x,y)$. Then as $S(2)$ excludes $D^{+}_{4}$ and $D^{-}_{4}$, we see that $\chi$ does not have a monochromatic copy of $\overrightarrow{C}_{3}$. ∎
@@
Let $\overrightarrow{C}_{4}$ denote the directed $4$-cycle. Then $\overrightarrow{C}_{4}\in\operatorname{Age}(S(3))$ (e.g. take all central angles to be $\frac{\pi}{2}$). Let $e\in S(3)$, let $\phi(x,y)\equiv xRy$, and let $\chi$ be the vertex-colouring externally defined by $e\in S(3)$ and $\phi(x,y)$. Then for any copy of $\overrightarrow{C}_{4}$ in $S(3)$, $e$ has an out-edge to one vertex of the copy and an in-edge from another vertex of the copy. Therefore $\chi$ has no monochromatic copy of $\overrightarrow{C}_{4}$. ∎
@@ 6.5. The rainbow graph, saturated version
$M$ has the pigeonhole property (see Definition 2.17): the proof is similar to that for the random graph. Thus we see that $M$ has vertex-Ramsey by Lemmas 2.18 and 2.19, and as $\operatorname{Th}(M)$ has QE, by Proposition 3.13 we have:
@@ 6.6. The rainbow graph, unsaturated version
Let =$\mathcal{L}=\{R_{i}:i<\omega\}$ be as in the previous example. We say that an $\mathcal{L}$-structure $A$ is a complete $\omega$-coloured graph if $A$ is an $\omega$-coloured graph (see Definition 6.10 and for all $a,b\in A$ with $a\neq b$, there exists $i<\omega$ with $aR_{i}b$.
@@ 6.7. Infinitely refining equivalence relations, saturated version
For $i<\omega$, let $\phi_{i}(x,y)\equiv xR_{i}y$. By Proposition 3.13, it suffices to check vertex-Ramsey for each $2$-colouring externally defined in an elementary extension by the formula $\phi_{i}(x,y)$ ($i<\omega$).
@@
To see this, we note that the saturated version of the structure $M$ with infinitely refining equivalence relations is isomorphic to the lexicographic product $M^{\prime}[N]$, where $M^{\prime}$ is the unsaturated version of the structure of infinitely refining equivalence relations described in the following section, and $N$ is a pure set in the language of equality. (See Section 7.) We will establish in Proposition 6.20 that $M^{\prime}$ is indivisible, and $N$ is clearly indivisible as well. We then see from the following fact that $M^{\prime}[N]$ is indivisible (since $M^{\prime}[N]$ is a reduct of $M^{\prime}[N]^{s}$), and thus in particular has vertex-Ramsey.
Fact 6.18 ([20, Proposition 2.21]).
@@ 6.9. Disjoint random graphs
Take the $\mathcal{L}$-formula $\sigma(x,y)\equiv(\exists\,z)(z\sim x\wedge z\sim y)$. Then the equivalence relation $s$ is $\varnothing$-definable in $M$ by $\sigma(x,y)$. As $\operatorname{Th}(M_{s})$ has QE, we have that $\operatorname{Th}(M)$ has “QE up to $\sigma(x,y)$”, i.e. each $\mathcal{L}$-formula is equivalent modulo $\operatorname{Th}(M)$ to a Boolean combination of $x\sim y$ and $\sigma(x,y)$. Therefore, by Lemma 3.8, it suffices to check externally definable Ramsey for $M$ for colourings externally defined by $x\sim y$ and by $\sigma(x,y)$. For a colouring $\chi:\binom{M}{a}\to 2$ externally defined by $e\in N\succeq M$ and $\phi(x,y)$, where $\phi(x,y)\equiv x\sim y$ or $\phi(x,y)=\sigma(x,y)$, we have that $\chi$ restricts to a colouring on $M_{0}$, and as $\operatorname{Age}(M)\subseteq\operatorname{Age}(M_{0})$ and $M_{0}$ has point-Ramsey, we are done. ∎
@@ 7. The lexicographic product
Let $\mathcal{L}^{s}$ be the expansion of $\mathcal{L}$ by a new binary relation symbol $s$, and let $M[N]^{s}$ denote the $\mathcal{L}^{s}$-structure resulting from expanding the $\mathcal{L}$-structure $M[N]$ by interpreting $s$ as: $((a_{1},b_{1}),(a_{2},b_{2}))\in s^{M[N]}\iff a_{1}=a_{2}$. Let $N^{s}$ denote the $\mathcal{L}^{s}$-structure resulting from expanding $N$ by the binary relation $s=N^{2}$. Then $M[N]^{s}=M[N^{s}]$.
@@
Proposition 7.2 ([20, Proposition 2.23]).
Let $M,N$ be $\mathcal{L}$-structures, and assume $M$ is transitive.
@@
Let $M,N$ be $\mathcal{L}$-structures, and assume $M$ is transitive. Let $A\subseteq M$ and $B\subseteq N$. Let $\bar{c},\bar{c}^{\prime}\in M$ and $\bar{d},\bar{d}^{\prime}\in N$. Then the following are equivalent:
  1. (1)
    $\operatorname{tp}(\overline{(c,d)}/A[B]^{s})=\operatorname{tp}(\overline{(c^{\prime},d^{\prime})}/A[B]^{s})$;
  2. (2)
    $\operatorname{tp}(\overline{c}/A)=\operatorname{tp}(\overline{c}^{\prime}/A)$ and $\operatorname{tp}(\overline{d}/B)=\operatorname{tp}(\overline{d}^{\prime}/B)$.
@@
1 $\Rightarrow$ 2 is clear. For the other direction, assume $\operatorname{tp}(\overline{c}/A)=\operatorname{tp}(\overline{c}^{\prime}/A)$ and $\operatorname{tp}(\overline{d}/A)=\operatorname{tp}(\overline{d}^{\prime}/A)$. Let $\phi(\overline{x},\overline{z})$ be an $\mathcal{L}^{s}$-formula such that $|\overline{z}|=|\overline{(c,d)}|$. By Proposition 7.2, there exist $k\in\mathbb{N}$ and $\mathcal{L}$-formulae $(\phi_{1}^{j}(\overline{x},\overline{z}))_{j<k},(\phi_{2}^{j}(\overline{x},\overline{z}))_{j<k}$ such that, for every pair of tuples $\overline{(\alpha,\beta)},\overline{(\gamma,\delta)}\in M[N]^{s}$ with $|\overline{(\alpha,\beta)}|=|\overline{x}|$ and $|\overline{(\gamma,\delta)}|=|\overline{z}|$, we have:
@@
$\iff$ $M[N]^{s}\models\phi(\overline{(a,b)},\overline{(c^{\prime},d^{\prime})})$
Lemma 7.4 ([20, Lemma 2.13]).
@@
Conversely, if $M$ is transitive, then for every $F\in\operatorname{Aut}(M[N]^{s})$, there exists an automorphism $f\in\operatorname{Aut}(M)$ and a family of automorphisms $(g_{a}\in\operatorname{Aut}(N):a\in M)$ such that for all $(x,y)\in M[N]^{s}$, $F(x,y)=(f(x),g_{x}(y))$.
In other words, if $M$ is transitive, then
@@
Lemma 7.6 ([21, Corollary 1.2.5]).
Let $M,M^{\prime},N,N^{\prime}$ be $\mathcal{L}$-structures. If $M\equiv M^{\prime}$ and $N\equiv N^{\prime}$, then $M[N]\equiv M^{\prime}[N^{\prime}]$.
@@
Let $M$ be transitive. If $M$ has FPTn, then so does $M[N]^{s}$.
@@
Let $p\in S_{n}(M[N]^{s})$. We realise $p$ as a tuple $\overline{(a,b)}$ in an elementary extension of $M[N]^{s}$, which, by Lemma 7.7, we may assume is $M^{\prime}[N^{\prime}]^{s}$ for some $M^{\prime}\succeq M$, $N^{\prime}\succeq N$.
Let $p_{1}=\operatorname{tp}_{M^{\prime}}(\overline{a}/M)$ and $p_{2}=\operatorname{tp}_{N^{\prime}}(\overline{b}/N)$. As $M$ has $FPT_{n}$, there is an $\operatorname{Aut}(M)$-invariant type $q_{1}\in\overline{\operatorname{Aut}(M)\cdot p_{1}}$. Let $\bar{c}\in M^{\prime\prime}\succeq M$ be a realisation of $q_{1}$, where $M^{\prime\prime}$ is some elementary extension of $M$. Since $M$ is transitive and $q_{1}$ is $\operatorname{Aut}(M)$-invariant, we have that $\bar{c}\in M^{\prime\prime}\setminus M$. Let $q=\operatorname{tp}_{M^{\prime\prime}[N^{\prime}]^{s}}(\overline{(c,b)}/M[N]^{s})$. By Lemma 7.3, $q$ depends only on $q_{1}$ and $p_{2}$, so we denote $q$ obtained as above from $q_{1}$ and $p_{2}$ by $q=:q_{1}[p_{2}]$. We claim that $q$ is $G$-invariant and lies in the $G$-orbit closure of $p$.
@@
By Lemma 7.4, the function $\sigma^{\prime\prime}_{2}:M^{\prime\prime}[N]^{s}\to M^{\prime\prime}[N]^{s}$ given by
$\sigma^{\prime\prime}_{2}(x,y)=\begin{cases}(x,g_{x}(y))&x\in M\\ (x,y)&x\not\in M\end{cases}$
is an automorphism of $M^{\prime\prime}[N]^{s}$. We have that $\sigma^{\prime\prime}_{2}$ extends $\sigma_{2}$, and using the fact that $\bar{c}\in M^{\prime\prime}\setminus M$, we have $\sigma^{\prime\prime}_{2}(\overline{(c,b)})=\overline{(c,b)}$. We therefore have that $q$ is $\sigma_{2}$-invariant.
To see that $q$ is $\sigma_{1}$-invariant, we recall that $q_{1}$ is $\operatorname{Aut}(M)$-invariant, so in particular is invariant under $f$. Therefore, $\sigma_{1}(q)=\sigma_{1}(q_{1}[p_{2}])=f(q_{1})[p_{2}]=q_{1}[p_{2}]=q$. We have now shown that $q$ is $\operatorname{Aut}(M)$-invariant.
@@
Since $q=q_{1}[p_{2}]$, by definition and by Lemma 7.3, there are $\mathcal{L}$-formulae $\phi_{1},\phi_{2}$, such that $\phi_{1}\in q_{1}$ and $\phi_{2}\in p_{2}$, such that whenever $q^{\prime}=q^{\prime}_{1}[q^{\prime}_{2}]$ with $\phi_{1}\in q^{\prime}_{1}$ and $\phi_{2}\in q^{\prime}_{2}$, we have that $\phi\in q$. Now, $q_{1}\in\overline{\operatorname{Aut}(M)\cdot p_{1}}$, and therefore there is some $g_{1}\in\operatorname{Aut}(M)$ such that $\phi_{1}\in g_{1}\cdot p_{1}$. Finally, let $g\in\operatorname{Aut}(M[N]^{s})$ be defined as $g(x,y):=(g_{1}(x),y)$. By Lemma 7.4, $g\in\operatorname{Aut}(M[N]^{s})$, and $g\cdot p=g\cdot p_{1}[p_{2}]=(g_{1}\cdot p)[p_{2}]\ni\phi$. ∎

### v4

@@ 6.3. Free amalgamation classes via Nešetřil-Rödl
Let $n\geq 2$, and let =$\mathcal{L}=\{R_{i}:i\in I\}$ with each $R_{i}$ of arity $n$. Let $M$ be a Fraïssé $\mathcal{L}$-structure with age $\mathcal{K}$ such that:
  • each $R_{i}^{M}$ is symmetric;
  • $\mathcal{K}$ has free amalgamation;
  • $\operatorname{Th}(M)$ has QE.
Then $M$ has the externally definable Ramsey property.
@@
(Similar to [15, Example 5.6].) Let $\mathcal{K}_{\leq}$ be the order expansion of $\mathcal{K}$ with Fraïssé limit $M_{\leq}$ (with $M$ the $\mathcal{L}$-reduct of $M_{\leq}$). By Theorem 6.5, we see that $M_{\leq}$ has the RP. To show that $M$ has the externally definable RP, by Proposition 3.13 it suffices to check the $\bar{a}$-RP for externally atomic-definable $2$-colourings, where $|\bar{a}|<n$. Let $i\in I$ and $\bar{a}\in\mathcal{K}$, $|\bar{a}|<n$. Let $\chi:\binom{M}{\bar{a}}\to 2$ be externally defined by $\bar{e}\in N\succeq M$ (where $|\bar{e}|+|\bar{a}|=n$) and $\phi(\bar{x},\bar{y})\equiv R_{i}(\bar{x},\bar{y})$. Let $B\in\mathcal{K}$.
@@ 6.4. The directed graphs $S(2)$ and $S(3)$
The dense local order $S(2)$ is defined as follows. Let =$\mathcal{L}=\{R\}$ with $R$ binary. Let $\mathbb{T}$ denote the unit circle in the complex plane, and let $S(2)$ be the $\mathcal{L}$-structure whose domain consists of the points of $\mathbb{T}$ with rational argument, and where $aRb$ iff the angle subtended at the origin by the directed arc from $a$ to $b$ is less than $\pi$. As the domain of $S(2)$ excludes antipodal points, $S(2)$ is a tournament.
We have that $S(2)$ is ultrahomogeneous - see [17] for further details.
@@
Take $e\in S(2)$, and let $\phi(x,y)\equiv xRy$. Let $\chi$ be the vertex-colouring of $S(2)$ externally defined by $e\in S(2)$ and $\phi(x,y)$. It is then immediate that $\chi$ does not have a monochromatic copy of $\overrightarrow{C}_{3}\in\operatorname{Age}(S(2))$. ∎
@@
Let $\overrightarrow{C}_{4}$ denote the directed $4$-cycle. Then it is straightforward to check that $\overrightarrow{C}_{4}\in\operatorname{Age}(S(3))$. Let $e\in S(3)$, let $\phi(x,y)\equiv xRy$, and let $\chi$ be the vertex-colouring externally defined by $e\in S(3)$ and $\phi(x,y)$. Then for any copy of $\overrightarrow{C}_{4}$ in $S(3)$, $e$ has an out-edge to one vertex of the copy and an in-edge from another vertex of the copy. Therefore $\chi$ has no monochromatic copy of $\overrightarrow{C}_{4}$. ∎
@@ 6.5. The rainbow graph, saturated version
$M$ has the pigeonhole property (see Definition 2.17): the proof is similar to that for the random graph. Thus we see that $M$ has the vertex-Ramsey property by Lemmas 2.18 and 2.19, and as $\operatorname{Th}(M)$ has QE, by Proposition 3.13 we have:
@@ 6.6. The rainbow graph, unsaturated version
Let =$\mathcal{L}=\{R_{i}:i<\omega\}$ be as in the previous example. We say that an $\mathcal{L}$-structure $A$ is a complete $\omega$-coloured graph if $A$ is an $\omega$-coloured graph (see Definition 6.10) and for all $a,b\in A$ with $a\neq b$, there exists $i<\omega$ with $aR_{i}b$.
@@ 6.7. Infinitely refining equivalence relations, saturated version
For $i<\omega$, let $\phi_{i}(x,y)\equiv xR_{i}y$. By Proposition 3.13, it suffices to check the vertex-Ramsey property for each $2$-colouring externally defined in an elementary extension by the formula $\phi_{i}(x,y)$ ($i<\omega$).
@@
To see this, we note that the saturated version of the structure $M$ with infinitely refining equivalence relations is isomorphic to the lexicographic product $M^{\prime}[N]$, where $M^{\prime}$ is the unsaturated version of the structure of infinitely refining equivalence relations described in the following section, and $N$ is a pure set in the language of equality. (See Section 7.) We will establish in Proposition 6.20 that $M^{\prime}$ is indivisible, and $N$ is clearly indivisible as well. We then see from the following fact that $M^{\prime}[N]$ is indivisible (since $M^{\prime}[N]$ is a reduct of $M^{\prime}[N]^{s}$), and thus in particular has the vertex-Ramsey property.
Fact 6.18 ([19, Proposition 2.21]).
@@ 6.9. Disjoint random graphs
Take the $\mathcal{L}$-formula $\sigma(x,y)\equiv(\exists\,z)(z\sim x\wedge z\sim y)$. Then the equivalence relation $s$ is $\varnothing$-definable in $M$ by $\sigma(x,y)$. As $\operatorname{Th}(M_{s})$ has QE, we have that $\operatorname{Th}(M)$ has “QE up to $\sigma(x,y)$”, i.e. each $\mathcal{L}$-formula is equivalent modulo $\operatorname{Th}(M)$ to a Boolean combination of $x\sim y$ and $\sigma(x,y)$. Therefore, by Lemma 3.8 and a straightforward adaptation of Lemma 3.12, it suffices to check the externally definable RP for $M$ for colourings externally defined by $x\sim y$ and by $\sigma(x,y)$. For a colouring $\chi:\binom{M}{a}\to 2$ externally defined by $e\in N\succeq M$ and $\phi(x,y)$, where $\phi(x,y)\equiv x\sim y$ or $\phi(x,y)=\sigma(x,y)$, we have that $\chi$ restricts to a colouring on $M_{0}$, and as $\operatorname{Age}(M)\subseteq\operatorname{Age}(M_{0})$ and $M_{0}$ has the point-Ramsey property, we are done. ∎
@@ 7. The lexicographic product
Let $\mathcal{L}^{s}$ be the expansion of $\mathcal{L}$ by a new binary relation symbol $s$, and let $M[N]^{s}$ denote the $\mathcal{L}^{s}$-structure which is the expansion of the $\mathcal{L}$-structure $M[N]$ by:
$M[N]^{s}\models s((a_{1},b_{1}),(a_{2},b_{2}))\iff a_{1}=a_{2}.$
Let $N^{s}$ denote the $\mathcal{L}^{s}$-structure resulting from expanding $N$ by the binary relation $s=N^{2}$. Then $M[N]^{s}=M[N^{s}]$.
@@
Proposition 7.2 ([19, Proposition 2.23]).
Let $M,N$ be $\mathcal{L}$-structures, and assume $\operatorname{Th}(M)$ is transitive (i.e. any two points of $M$ are isomorphic as substructures).
@@
Let $M,N$ be $\mathcal{L}$-structures, and assume $\operatorname{Th}(M)$ is transitive. Let $A\subseteq M$ and $B\subseteq N$ with $A,B\neq\varnothing$. Let $\bar{c},\bar{c}^{\prime}\in M$ and $\bar{d},\bar{d}^{\prime}\in N$. Suppose that $\operatorname{tp}(\overline{c}/A)=\operatorname{tp}(\overline{c}^{\prime}/A)$ and $\operatorname{tp}(\overline{d}/B)=\operatorname{tp}(\overline{d}^{\prime}/B)$. Then $\operatorname{tp}(\overline{(c,d)}/A[B]^{s})=\operatorname{tp}(\overline{(c^{\prime},d^{\prime})}/A[B]^{s})$.
@@
Let $\phi(\overline{x},\overline{z})$ be an $\mathcal{L}^{s}$-formula such that $|\overline{z}|=|\overline{(c,d)}|$. By Proposition 7.2, there exist $k\in\mathbb{N}$ and $\mathcal{L}$-formulae $(\phi_{1}^{j}(\overline{x},\overline{z}))_{j<k},(\phi_{2}^{j}(\overline{x},\overline{z}))_{j<k}$ such that, for every pair of tuples $\overline{(\alpha,\beta)},\overline{(\gamma,\delta)}\in M[N]^{s}$ with $|\overline{(\alpha,\beta)}|=|\overline{x}|$ and $|\overline{(\gamma,\delta)}|=|\overline{z}|$, we have:
@@
$\iff$ $M[N]^{s}\models\phi(\overline{(a,b)},\overline{(c^{\prime},d^{\prime})}).$
Lemma 7.4 ([19, Lemma 2.13]).
@@
Conversely, if $\operatorname{Th}(M)$ is transitive, then for every $F\in\operatorname{Aut}(M[N]^{s})$, there exists an automorphism $f\in\operatorname{Aut}(M)$ and a family of automorphisms $(g_{a}\in\operatorname{Aut}(N):a\in M)$ such that for all $(x,y)\in M[N]^{s}$, $F(x,y)=(f(x),g_{x}(y))$.
In other words, if $\operatorname{Th}(M)$ is transitive, then
@@
Lemma 7.6 ([20, Corollary 1.2.5]).
Let $M,M^{\prime},N,N^{\prime}$ be $\mathcal{L}$-structures. If $M\equiv M^{\prime}$ and $N\equiv N^{\prime}$, then $M[N]^{s}\equiv M^{\prime}[N^{\prime}]^{s}$.
@@
Let $\operatorname{Th}(M)$ be transitive. If $M$ has FPTn, then so does $M[N]^{s}$.
@@
Let $p\in S_{n}(M[N]^{s})$. We realise $p$ as a tuple $\overline{(a,b)}$ in an elementary extension of $M[N]^{s}$, which, by Lemma 7.7, we may assume is $M^{\prime}[N^{\prime}]^{s}$ for some $M^{\prime}\succeq M$, $N^{\prime}\succeq N$. We will also assume that $N^{\prime}$ is strongly $(|N|+\aleph_{0})^{+}$-homogeneous.
Let $p_{1}=\operatorname{tp}_{M^{\prime}}(\overline{a}/M)$ and $p_{2}=\operatorname{tp}_{N^{\prime}}(\overline{b}/N)$. As $M$ has $FPT_{n}$, there is an $\operatorname{Aut}(M)$-invariant type $q_{1}\in\overline{\operatorname{Aut}(M)\cdot p_{1}}$. Let $\bar{c}\in M^{\prime\prime}\succeq M$ be a realisation of $q_{1}$, where $M^{\prime\prime}$ is some elementary extension of $M$. Since $\operatorname{Th}(M)$ is transitive and $q_{1}$ is $\operatorname{Aut}(M)$-invariant, we have that $\bar{c}\in M^{\prime\prime}\setminus M$. Let $q=\operatorname{tp}_{M^{\prime\prime}[N^{\prime}]^{s}}(\overline{(c,b)}/M[N]^{s})$. By Lemma 7.3, $q$ depends only on $q_{1}$ and $p_{2}$, so we denote $q$ obtained as above from $q_{1}$ and $p_{2}$ by $q=:q_{1}[p_{2}]$. We claim that $q$ is $G$-invariant and lies in the $G$-orbit closure of $p$.
@@
Using the fact that $N^{\prime}$ is strongly $(|N|+\aleph_{0})^{+}$-homogeneous, extend each $g_{m}$ (for $m\in M$) to an automorphism $g^{\prime}_{m}$ of $N^{\prime}$. By Lemma 7.4, the function $\sigma^{\prime\prime}_{2}:M^{\prime\prime}[N^{\prime}]^{s}\to M^{\prime\prime}[N^{\prime}]^{s}$ given by
$\sigma^{\prime\prime}_{2}(x,y)=\begin{cases}(x,g^{\prime}_{x}(y))&x\in M\\ (x,y)&x\not\in M\end{cases}$
is an automorphism of $M^{\prime\prime}[N^{\prime}]^{s}$. We have that $\sigma^{\prime\prime}_{2}$ extends $\sigma_{2}$, and using the fact that $\bar{c}\in M^{\prime\prime}\setminus M$, we have $\sigma^{\prime\prime}_{2}(\overline{(c,b)})=\overline{(c,b)}$. We therefore have that $q$ is $\sigma_{2}$-invariant.
To see that $q$ is $\sigma_{1}$-invariant, we recall that $q_{1}$ is $\operatorname{Aut}(M)$-invariant, so in particular is invariant under $f$. Therefore, $\sigma_{1}(q)=\sigma_{1}(q_{1}[p_{2}])=f(q_{1})[p_{2}]=q_{1}[p_{2}]=q$. This completes the proof of the claim that $q$ is $G$-invariant.
@@
Since $q=q_{1}[p_{2}]$, by definition and by Lemma 7.3, there are $\mathcal{L}$-formulae $\phi_{1},\phi_{2}$ with $\phi_{1}\in q_{1}$ and $\phi_{2}\in p_{2}$ such that whenever $q^{\prime}=q^{\prime}_{1}[q^{\prime}_{2}]$ with $\phi_{1}\in q^{\prime}_{1}$ and $\phi_{2}\in q^{\prime}_{2}$, we have that $\phi\in q$. Now, $q_{1}\in\overline{\operatorname{Aut}(M)\cdot p_{1}}$, and therefore there is some $g_{1}\in\operatorname{Aut}(M)$ such that $\phi_{1}\in g_{1}\cdot p_{1}$. Finally, let $g\in\operatorname{Aut}(M[N]^{s})$ be defined by $g(x,y):=(g_{1}(x),y)$. By Lemma 7.4, we have $g\in\operatorname{Aut}(M[N]^{s})$, and $g\cdot p=g\cdot p_{1}[p_{2}]=(g_{1}\cdot p_{1})[p_{2}]\ni\phi$. ∎